OpenStudy (anonymous):

Solve for x: |x + 2| + 16 = 14 Please Help Having Trouble.. :(

OpenStudy (anonymous):

-4: you subtract 16 and get (x+2)= -2 then you subtract 2 and get x=-4

OpenStudy (phi):

after you simplify to get | x+2 | = -2 (do you see how?) you now have a problem. the left hand side will be 0 or positive (because of the absolute value sign) in other words, you can never get the left side to be -2 (or any negative number) so no sol'n

OpenStudy (anonymous):

????? I'm still confused on all for this, Math isn't my strong subject,

OpenStudy (phi):

which part? if we start with |x + 2| + 16 = 14 do you know how to "add -16" to both sides?

OpenStudy (anonymous):

yes and no, its pretty confusing to me its the absolute value that throws me off

OpenStudy (phi):

write in -16 on both sides (don't both to simplify)

OpenStudy (phi):

and post it

OpenStudy (anonymous):

Okay

OpenStudy (phi):

you should get |x + 2| + 16 + -16 = 14 + -16

OpenStudy (phi):

now simplify the right-hand side (of the =)

OpenStudy (anonymous):

Okay and you get |x+2|=-2?

OpenStudy (phi):

yes

OpenStudy (phi):

now about the absolute value signs. they "do something" to the number inside them: if the number inside is positive, it leaves the number alone, but if the number is negative, it makes it positive.

OpenStudy (phi):

example: |1| becomes 1 |-1 | becomes 1 ok so far ?

OpenStudy (anonymous):

Okay then you subtract -2 from both sides Right?

OpenStudy (phi):

no, not yet. we have to get the idea clear. do you understand this part: the absolute value sign makes whatever is inside it positive.

OpenStudy (anonymous):

Yes

OpenStudy (phi):

think about this question: what number can you put inside the absolute value sign so that we get a negative number out? any guess?

OpenStudy (phi):

try to remember the absolute value sign makes whatever is inside it positive. in other words, | any number you can think of | will be made positive because that is what the | | do.

OpenStudy (phi):

if you had |x| = -2 what x would make this true? we could try numbers. Let's try x=-2 | -2 | = -2 ? the left side turns into +2 (that is what the | | does) 2 = -2? no. if we think about it, if the left side is positive for *any* number, then it will never be equal to a -2. in other words, there is no solution.